### v1

# CLoMo: Counterfactual Logical Modification with Large Language Models

Abstract
In this study, we delve into the realm of counterfactual reasoning capabilities of large language models (LLMs). Our primary objective is to cultivate the counterfactual thought processes within LLMs and rigorously assess these processes for their validity. Specifically, we introduce a novel task, Counterfactual Logical Modification (CLoMo), and a high-quality human-annotated benchmark. In this task, LLMs must adeptly alter a given argumentative text to uphold a predetermined logical relationship. To effectively evaluate a generation model’s counterfactual capabilities, we propose an innovative evaluation metric, the Logic-Aware Counterfactual Score to directly evaluate the natural language output of LLMs instead of modeling the task as a multiple-choice problem. Analysis shows that the proposed automatic metric aligns well with human preference. Our experimental results show that while LLMs demonstrate a notable capacity for logical counterfactual thinking, there remains a discernible gap between their current abilities and human performance.
1 Introduction
Despite large language models Arkoudas (2023); OpenAI (2022) perform strikingly in plenty of reasoning benchmarks Cobbe et al. (2021); Hendrycks et al. (2021a), late studies observe internal inconsistency in their reasoning processes Saparov and He (2023); Arkoudas (2023). The inconsistency is attributed to misunderstanding and misapplication of logical relations. However, logical relations in complex language reasoning are not yet properly quantified and evaluated.
Figure 1: Demonstration of CLoMo. An LLM is given an argument and two premises. The LLM needs to modify the statements in Argument such that the logical relation R switch to stand in state 2 instead of state 1.
Current studies on evaluating model reasoning are limited in both form and content. On the one hand, benchmarking complex reasoning is generally applying discrimination tasks such as multiple-choice questions Huang et al. (2023b); Hendrycks et al. (2021a); Chen et al. (2023); Suzgun et al. (2023), where accuracy and pass rate serve as the main evaluation metric. However, such evaluations over-simplify the goal of uncovering essential and subtle pitfalls in complex reasoning. For example, the reasoning processes could contain misconceptions in logical relations but give correct answers due to the data distribution Elazar et al. (2021); Saparov and He (2023). Therefore, evaluating the generated content would provide a more realistic measurement of model reasoning. On the other hand, unlike widely studied reasoning tasks such as math reasoning Cobbe et al. (2021); Hendrycks et al. (2021b) and standard exams OpenAI (2023); Huang et al. (2023b), counterfactual reasoning Starr (2022) as a fundamental evaluation of logical relations is less explored in the context of large language models. Previous literature studies counterfactual reasoning either in a multiple-choice manner Tandon et al. (2019); Qin et al. (2021) or applying labored human study to evaluate counterfactual generation Qin et al. (2019), leaving an effective evaluation of counterfactual generation unexplored.
In our study, we delve into the realm of evaluating large language models’ (LLMs) ability to generate counterfactually coherent thoughts. Specifically, we proposed an innovative evaluation system that quantitatively measures the evolution of information in statement pairs, ensuring that they adhere to a specified logical relationship. Our approach includes designing a specialized task where models are presented with mismatched argument-premise pairs bound by a specific logical relation. The objective for these models is to adeptly modify the argument text until the specified logical relation is satisfactorily established. In conjunction with this task, we have created the first dataset of its kind, comprising dual argument-premise pairs, each annotated with a defined logical relation. This dataset is vital for facilitating logically restricted counterfactual modifications, and we have enriched it with human-written modifications to serve as a benchmark for evaluation.
Our experimental investigations encompass a range of large language models, including GPT-4 OpenAI (2023) and GPT-3.5-Turbo OpenAI (2022), as well as smaller models from the LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b) families. Through these experiments, we have discerned that the task of CLoMo poses a significant challenge. It becomes evident that these models’ current counterfactual logical reasoning capabilities fall short of the desired proficiency. This observation underscores the need for further advancements in enhancing the counterfactual reasoning abilities of existing language models, paving the way for more sophisticated and logically coherent AI systems.
We summarize the contributions of this paper as follows:
1. We propose the task of Counterfactual Logical Modification, and contribute a corresponding CLoMo. to evaluate the counterfactual reasoning capability of LLMs in the scenario of complicated textual logical reasoning.
2. We propose the Logic-Aware Counterfactual Score for the logically consistent generation of large language models.
3. We conduct experiments on representative LLMs (GPT-3.5, GPT-4) and smaller language models (the LLaMA and LLaMA 2 families) and find that CLoMo is a very challenging task and the counterfactual logical reasoning ability of the existing model needs to be improved.
R State 1 State 2
Necessary Assumption (38.5%)
Argument: Statement1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions backed by weak evidence, politicians should not generally be criticized for ignoring social science, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
Premise1: The failure of scientists to agree that a claim within their specialty is true can indicate that the evidence for the claim is not strong.
${}^{\prime}$: Statement 1: Journalist: the advice of social scientists is frequently overlooked by politicians making social policy. Statement 2: Because it is not unreasonable to discount scientific assertions, politicians should not generally be criticized for ignoring social science unless social scientists agree on the same claim, for social scientists, unlike physical scientists, seldom agree on the claims made even within their own specialty.
Premise2: Politicians should follow the advice of experts on issues about which those experts agree among themselves.
Sufficient Assumption (6.6%)
Argument: Statement 1: Caffeine can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of many diverse species of plants is not accidental but evolved as a defense for those plants.
Premise1: Caffeine-producing plants or their ancestors have sometimes been fed upon by creatures sensitive to caffeine.
${}^{\prime}$: Statement 1: Caffeine produced for plant species’ own defense can kill or inhibit the growth of the larvae of several species of insects. Statement 2: One recent experiment showed that tobacco hornworm larvae die when they ingest a preparation that consists, in part, of finely powdered tea leaves, which contain caffeine. Statement 3: This result is evidence for the hypothesis that the presence of non-negligible quantities of caffeine in various parts of tobacco plant is not accidental but evolved as a defense for it.
Premise2: The tobacco plant is among the plant species that produce caffeine for their own defense.
Strengthen (18.7%)
Argument: Statement1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But many expenses do not occur monthly ; taking into account only monthly expenses can cause a business to overexpand. Statement2: So the use of a cash-flow statement is critical for all businesses.
Premise1: A cash-flow statement is the only way to track both monthly expenses and expenses that are not monthly.
${}^{\prime}$: Statement 1: In contemplating major purchases, businesses often consider only whether there is enough money left from monthly revenues after paying monthly expenses to cover the cost of the purchase. But there are many expenses every month ; taking into account these expenses incorrectly can cause a business to overexpand. Statement 2: So the use of a cash-flow statement is critical for all businesses.
Premise2: Only a cash-flow statement can accurately document all monthly expenses.
Weaken (36.2%)
Argument: Statement1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat, and it takes several pounds of grain to produce one pound of meat. Statement2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic grain production are highly unlikely, baurisia is soon likely to become an importer of grain.
Premise1: It is more economical for baurisians to import meat than grain.
${}^{\prime}$: Statement 1: The country of baurisia has, until now, been self-sufficient in both grain and meat. However, with growing prosperity in baurisia has come a steadily increasing per capita consumption of meat. Statement 2: Therefore, since per capita income in baurisia is almost certain to rise further but increases in domestic meat production are highly unlikely, baurisia is soon likely to become an importer of meat.
Premise2: The per capita consumption of meat in baurisia is roughly the same across all income levels.
Table 1: Example questions from the CLoMo benchmark, with the proportion of each logical relation.
2 Related Works
From Complex Reasoning to Counterfactual Reasoning
Complex reasoning has been highly concerned as a significant yet challenging task for inspecting advanced artificial intelligence. For example, for solving commonsense reasoning problems Talmor et al. (2019, 2021); Huang et al. (2019); Bhagavatula et al. (2020); Sap et al. (2019), the models Yasunaga et al. (2021, 2022); Liu et al. (2021); Huang et al. (2021b) are required to reasonable applying commonsense knowledge to conduct the reasoning process for the final answer. Furthermore, multi-step reasoning needs the models to perform multiple reasoning steps while maintaining consistency and faithfulness. To achieve this, synthetic compositional reasoning tasks Betz (2020); Tafjord et al. (2021); Han et al. (2022); Saparov and He (2023) incorporates first-order logic to inspect and improve models logical consistency Pan et al. (2023); Olausson et al. (2023); Sanyal et al. (2022); Yang and Deng (2021). Moreover, real-scenario compositional reasoning Yu et al. (2020); Liu et al. (2020); Dalvi et al. (2021); Huang et al. (2022) joins commonsense, consider the uncertainty of events in multi-step logical reasoning, which challenge current models Bao et al. (2023); Xu et al. (2023); Jiao et al. (2023, 2022); Huang et al. (2023a, 2021a) to solve real-world reasoning problems with faithfulness. Additionally, the more faithful models should be able to consider counterfactuals. For example, answering questions given counterfactual conditions Yu et al. (2023); Tandon et al. (2019); Qin et al. (2021), or narrating a counterfactual scenario Qin et al. (2019). However, previous studies on counterfactual reasoning barely pay attention to the logical consistency or faithfulness of models. Therefore, in this work, we propose the task of Counterfactual Logical Modification that challenges the model to satisfy a given logical relation restriction while generating counterfactuals.
Evaluation of LLM Reasoning
Reasoning poses a great challenge to LLMs. Currently, there is an increasing interest in the reasoning ability of LLMs. Evaluations are carried out from several perspectives, including mathematical reasoning, common-sense reasoning, logical reasoning, and domain-specific reasoning Chang et al. (2023); Zhong et al. (2023); Bang et al. (2023); Liu et al. (2023). However, most current methods of reasoning evaluation focus primarily on the accuracy of the final answer and neglect a comprehensive assessment of the reasoning process of LLMs. Such evaluation is not ideal for understanding the reasoning ability of models, as it ignores situations where models may obtain correct answers through unfaithful or spurious reasoning shortcuts Saparov and He (2023). Some recent research has started to evaluate the validity of the intermediate reasoning steps of LLMs Golovneva et al. (2022); Prasad et al. (2023). However, they mainly focus on the relationship between the intermediate step and the final answer rather than measuring whether the model understands the intermediate reasoning process. This paper proposes a novel logical reasoning benchmark that requires intermediate counterfactual thinking under logical relation restrictions. This leads to a more in-depth study of the model’s intermediate reasoning process.
3 CLoMo Benchmark
3.1 Task Definition
The desideratum is to harvest LLM counterfactual thinking and then investigate the validation of the thinking and its alignment with human counterfactual thinking. To achieve this, the LLM should generate its counterfactual thinking under proper logical scenarios. We design a task of counterfactual modification of argument text given a perturbation of premise given a static logical relation.
A demonstration of the proposed Counterfactual Logical Modification is shown in Figure 1. An LLM is given the instruction as shown on the left-hand side, which can be illustrated by the diagram on the right-hand side. In the given instruction, Argument and Premise1 are related by a logical relation. We consider four main relations in practice, which are (R1) the premise provides a necessary assumption to the argument, (R2) the premise provides a sufficient assumption to the argument, (R3) the premise strengthen the argument, and (R4) the premise weaken the argument. The Argument and Premise1then constitute State 1 of the logical relation R. The additional Premise2 perturbs the logical relation R. The goal for the LLM is to maintain a State 2 with the given Premise2 and a modified ${}^{\prime}$ that R stands. To this end, it should properly edit the statements in Argument until the goal is reached.
3.2 Benchmark Construction
Since applicable data for the proposed task is lacking, we build a benchmark dataset by carefully selecting argument texts and collecting human annotation of the modified ${}^{\prime}$. We choose to use ReClor Yu et al. (2020) as the source data, considering that ReClor includes standardized multiple-choice questions on argument texts and logical relations from LSAT. We then recruit domain experts to conduct the annotation.
Given a ReClor question with context (the argument text), question, options, and the correct answer option, an annotator is required to provide a chosen wrong option (as the Premise2) and a corresponding modified context (i.e., the modified ${}^{\prime}$) to form a data point. The annotator is first instructed to read the whole question and comprehend the in-line logical relations, then choose one of the wrong options. After that, he/she edits the context by deleting, adding, or replacing text spans in the statements. The number of editions and the length of edited text spans are unrestricted as long as the statement partition is maintained.
We then post-process the question and the annotation so that for each data point, Argumnet comes from the original context, Premise1 comes from the correct option, and Premise2 and ${}^{\prime}$ come from the annotation.
4 SES: Self-Evaluation Scores
We aim to use the proven complex reasoning capabilities of the large language model itself to perform faster and more efficient reasoning evaluations of complex reasoning tasks that do not have easy access to standard/human-tested answers. Specifically, we split the scenario of complex logical reasoning evaluation into several discriminative tasks that LLMs have already seen and have been heavily trained on through logical conceptual graphs in order for the model to perform high-precision reasoning evaluation. We then collect the evaluations of these simple tasks and compute the ratings of the complex reasoning tasks based on the structure of the logical concept graph.
Figure 2: The concept graph of counterfactual logical modification.
$c(r|p_{1},a)$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
Argument: <$a$>
Premise: <$p_{1}$>
$c(r|p_{2},a^{\prime})$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
Argument: <$a^{\prime}$>
Premise: <$p_{2}$>
$c(r|p_{2},a)$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
Argument: <$a$>
Premise: <$p_{2}$>
Table 2: Prompts for LLM decomposed evaluation task.
Counterfactual Modification Concept Graph Figure 2 To make the logic of counterfactual reasoning hold, we have a pair of primitive states $p_{1}$ (Premise1) and $a$ (Argument) from $p_{1}$ and $a$ satisfying the relation $r$, i.e., $\text{Pr}(r|p_{1},a)$ approaches $1$. The other claim $p_{2}$ (Premise2) and the modified ($a^{\prime}$) ${}^{\prime}$satisfy the same relation $r$, i.e., $\text{Pr}(r|p_{2},a^{\prime})$ approaches $1$. In contrast, the relation between $p_{2}$ and the Argumentshould not satisfy the relation $r$, i.e., $\text{Pr}(r|p_{2},a)$ approaches 00.
Decomposed Self-Evaluation Metric As the demonstrated complex logical reasoning ability of current large language models, we let the large language model estimate $\text{Pr}(r|p_{1},a)$, $\text{Pr}(r|p_{2},a^{\prime})$, and $\text{Pr}(r|p_{2},a)$, respectively. Specifically, we design a few binary classification tasks for the large language models, so that the probabilities are simplified to $c(r|p_{1},a)$, $c(r|p_{2},a^{\prime})$, and $c(r|p_{2},a)$$\in\{0,1\}$. Finally, the overall logical modification score is computed by:
$s=c(r|p_{1},a)\times c(r|p_{2},a^{\prime})-c(r|p_{2},a)\times c(r|p_{2},a^{\prime})$ (1)
Alignment with Human Evaluation We randomly selected 50 samples from the test set and examined how well the self-evaluation score (SES) matches human evaluation. Specifically, we use GPT-4 to implement SES. We recruit experts in argumentation to evaluate the modified arguments generated by GPT-4 on the selected 50 samples, scoring $1$ for good answers and 00 for bad answers. The Cohen’s Kappa coefficient between the human annotators is $\kappa=0.6785$, indicating substantial consistency of human perspective. We then use the self-evaluation score to evaluate the same group of modified arguments again. Cohen’s kappa coefficient between human and self-evaluation score is ${\kappa}=0.4391$. This indicates that the self-evaluation score is consistent with human evaluation. Therefore, we can use the self-evaluation score for more efficient logical modification evaluation. Moreover, since we can always check the logical reasoning ability of the model and its alignment with human evaluation, we can use a different model as the backbone to calculate this score.
SES NA SA S W
GPT-3.5-Turbo OpenAI (2022) 0.335 0.405 0.143 0.486 0.222
GPT-4 OpenAI (2023) 0.475 0.544 0.643 0.714 0.250
Human Performance 0.580 0.456 0.500 0.771 0.639
Table 3: Self-evaluation scores of LLM-generated counterfactual statements. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
Few
Zero
CoT
Inference-Only LLaMA-7b
0.340
0.435
0.440
LLaMA-13b
0.435
0.410
0.425
LLaMA 2-7b
0.270
0.145
0.115
LLaMA 2-13b
0.300
0.045
0.430
LoRA LLaMA-7b
0.260
0.285
0.295
LLaMA-13b
0.415
0.420
0.395
LLaMA 2-7b
0.085
0.135
0.145
LLaMA 2-13b
0.435
0.305
0.300
Fine-Tuning LLaMA-7b
0.385
0.445
0.085
LLaMA-13b
0.494
0.291
0.430
LLaMA 2-7b
0.633
0.557
0.494
LLaMA 2-13b
0.506
0.468
0.494
Table 4: SES results of LLaMA and LLaMA 2 families. Few: few-shot demonstrations. Zero: zero-shot demonstrations. CoT: chain-of-thought prompting.
Figure 3: Per-relation performances of LLaMA and LLaMA 2 families.
5 Experiments
5.1 Main Results
We first evaluate the large language models. We evaluate GPT-3.5-Turbo and GPT-4, respectively. At the same time, we recruited experts in linguistics to annotate the answers with humans. Table 3 demonstrates the results.
In the results, we see that the overall performance of the human experts on CLoMois 0.542, indicating that CLoMois quite a challenging task. In terms of the performance of the 4 logical relations, strengthen and weaken are relatively easy, while necessary assumption and sufficient assumption are more challenging for humans. In addition, the performance of GPT-4 is slightly lower than that of humans, but it also demonstrates strong counterfactual logical reasoning ability. And GPT-3.5-Turbo performs inferior to both human and GPT-4. Overall, large language models show great potential for counterfactual logical reasoning.
5.2 Results on Smaller Models
We further evaluate LLaMA and LLaMA 2 on CLoMo. We conduct fine-tuning with LoRA and full fine-tuning. We also directly evaluate the pre-trained models. We use three formats of training and testing data points, including few-shot demonstrations (Few), zero-shot demonstrations (Zero), and chain-of-thought prompting (CoT). The evaluation results are shown in Table 4.
In general, the pre-trained LLaMA and LLaMA 2 models achieve a certain level of counterfactual reasoning, and model scale is beneficial, but still fall short of large language models such as GPT-4. Fine-tuning with LoRA does not improve the performances. Full fine-tuning is more beneficial than fine-tuning with LoRA for this task. On the other hand, few-shot demonstrations in the data points help solve the task better. This suggests that the models can learn from imitation to solve this challenging task. Also, chain-of-thought prompting helps the counterfactual reasoning of the model to some extent.
In conclusion, smaller models in counterfactual logical reasoning still need to be improved.
5.3 Impacts of Logical Relations
Among the four logical relations, many of the models can not predict the Sufficient Assumption. Especially with zero-shot demonstration data. For example, both LLaMA 2-7b and LLaMA 2-13b fine-tuned with LoRA gets SES=0 in this logical relation. Therefore, Sufficient Assumption seems to be the hardest relation for the models. In contrast, the models generally perform better in Necessary Assumptions. Some models work better after fine-tuning. For example, LLaMA-7b improves from 0.4 to 0.429 after fine-tuned with LoRA and to 0.514 after full fine-tuning, and improves from 0.086 to 0.486 after fine-tuned with LoRA and 0.500 after full fine-tuning. While in some settings, for example, using the chain-of-thought prompting data, LLaMA-7b performs well with the pre-trained parameters, and fine-tuning with LoRA or full parameters both induce performance drop.
In conclusion, the extent to which the model understands different logical relations varies considerably. Therefore, an improved understanding of all logical relations is needed to improve models’ counterfactual logical reasoning ability.
Figure 4: A successful case of counterfactual modification generated by GPT-4, which makes a logically consistent ${}^{\prime}$ and is accordant with human reasoning.
Figure 5: An inferior case of counterfactual modification generated by GPT-4. GPT-4 modifies Statement 1 by replacing intellectual activities with physical activities, and the logical relation restriction is not satisfied.
5.4 Case Study
We compare the modification by GPT-4 and humans.
Figure 4 shows a case about sufficient assumption. That is, whether premise is a sufficient assumption for argument. The focus of the two premises here is to predict the impact of the charter provision. Switching from Premise 1 to Premise 2, the focus of discussion changes from the group of major powers to other nations in response to threats to world peace. This mainly affects the elaboration of Statement 3. GPT-4 has revised Statement 3 accordingly. Humans also made changes to Statement 3. Both revisions emphasized the influence of the five major powers in Statement 3, which corresponded to Premise 2. We find that GPT-4 can handle complex logical reasoning and counterfactual reasoning to a certain extent.
Figure 5 is about weakening an argument. The argument is on human intellectual development. Statement 2 provides evidence to support Statement 1. Premise 1 is on medical conditions and treatments providing counterexamples, and Premise 2 is on inaccuracies in research data. Human modifies Statement 2 to emphasize conversation, thus satisfying the logical conflict with the inaccuracies in research data described in Premise 2, thus satisfying the weaken relation. GPT-4 modifies Statement 1 by replacing intellectual activities with physical activities such as sports or gym. Intellectual/physical activities have some counterfactual contrast, but in the context of the argument at hand, the logical relation restriction is not satisfied. We find that GPT-4 can still be flawed in counterfactual reasoning.
In conclusion, complex counterfactual reasoning is challenging for large language models and needs improvements.
6 Conclusion
In this paper, we study large language models’ counterfactual reasoning capability under the constraint of proper logical relations. Our goal is to harvest LLM counterfactual thinking and then investigate the validation of the thinking and its alignment with human counterfactual thinking. To this end, we introduce a novel task, Counterfactual Logical Modification, that requires the LLMs to conduct counterfactual modification with logical restriction, where LLMs need to appropriately modify an argument text so that a specified logical relation stands. This task not only challenges the LLMs but also lays the groundwork for more nuanced and logically coherent counterfactual reasoning. To ensure a comprehensive evaluation, we then construct a benchmark dataset CLoMo by gathering human annotations and constructed a tailored dataset specifically designed for this purpose. Moreover, to evaluate current intelligent LLMs, we propose a logic-aware counterfactual score that leverages LLMs’ logical reasoning foundations. Crucially, we have aligned this logic-aware counterfactual score with human evaluations to ensure its relevance and accuracy in assessing the LLMs’ performance. Furthermore, we evaluate smaller language models including the LLaMA and LLaMA 2 families on CLoMo. The experimental results find that there is still room for LLMs to develop their counterfactual logical thinking. This gap underscores the need for further refinement and development in LLMs’ counterfactual reasoning faculties. Our findings thus open new pathways for future research to bridge this gap and enhance counterfactual reasoning in large language models.